OpenStudy (anonymous):

you must panel 10 rooms in an apartment building. Each room is 12' by 16' with 8' ceilings. Paneling sheets are 4' by 8' assume any paneling cut is waste. How many panels are needed to panel all of the rooms?

OpenStudy (anonymous):

area of 4 walls in one room = 12 * 16 * 4 total area of 10 rooms = 12 * 16 * 4 *10 area of one panel = 4 * 8 number of panels = 12 * 16 * 4 * 10 ------------- = 240 4 * 8

OpenStudy (anonymous):

final asnwer is 240? thanks

OpenStudy (anonymous):

yes 240 panels

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ur welcome